### v2

@@ Abstract
We study the use of large language model-based agents for interacting with software via web browsers. Unlike prior work, we focus on measuring the agents’ ability to perform tasks that span the typical daily work of knowledge workers utilizing enterprise software systems. To this end, we propose WorkArena, a remote-hosted benchmark of 29 tasks based on the widely-used ServiceNow platform. We also introduce BrowserGym, an environment for the design and evaluation of such agents, offering a rich set of actions as well as multimodal observations. Our empirical evaluation reveals that while current agents show promise on WorkArena, there remains a considerable gap towards achieving full task automation. Notably, our analysis uncovers a significant performance disparity between open and closed-source LLMs, highlighting a critical area for future exploration and development in the field.
@@
Figure 1: Overview of contributions: (a) WorkArena is a benchmark of 29 web tasks and 18,050 unique instances that cover common ways of interacting with the ServiceNow Platform, a widely-used enterprise software platform. (b) BrowserGym is a Python environment for designing and evaluating web agents, which includes a rich set of actions and multimodal observations (shown here the HTML contents of the page, its accessibility tree, and the raw pixels after browser rendering).
@@ 1 Introduction
  • WorkArena: A realistic benchmark of enterprise-related tasks for web agents comprising 18,050 unique task instances (§ 3, Fig. 1a);
  • BrowserGym: A new framework for the development and evaluation of web agents, compatible with previous benchmarks like WebArena (Zhou et al., 2023), MiniWoB (Liu et al., 2018; Shi et al., 2017a) and WebShop (Yao et al., 2022), that offers a richer set of multimodal observations (e.g., screenshot, accessibility tree, screen coordinates), a broader set of actions (e.g., Python code and high-level primitives), and supports chat-based interactions (§ 4, Fig. 1b); Surprisingly, these features contribute to bringing our GPT-4 agent at the top of the leaderboard on WebArena, with a score of 25.4%, contrasting with the score of the original paper, 14.4%.
@@ 2 Related Works
Benchmarks for web agents:  Early benchmarks for web agents were based on synthetic web environments where agents were tasked with performing low-level keyboard and mouse actions (Shi et al., 2017b). Notable examples are MiniWoB (Shi et al., 2017a; Liu et al., 2018), which offer a collection of 125 toy web tasks ranging from clicking a specific button to using a basic text editor, and WebShop (Yao et al., 2022), a simulated e-commerce website with shopping tasks that require searching and browsing a catalog of items. More recently, Zhou et al. (2023) introduced WebArena, a collection of 190 tasks based on realistic websites that emulate real-world domains such as e-commerce, social forums, collaborative software development, and content management. WebArena is a notoriously challenging benchmark, with a success rate of 14% for a state-of-the-art web agent based on GPT-4, and 78% for human agents. Deng et al. (2023) proposed Min2Web, a large-scale dataset of 2,000 web interactions from 137 websites curated by human annotators. Similarly, Lù et al. (2024) propose WebLINX, a curated dataset of web interactions composed of 2337 expert demonstrations from 155 different real-world websites. In WebLINX, each task is composed of a turn-based chat dialogue averaging 43 interactions per task. Last, He et al. (2024) propose 300 information-retrieval tasks, from 15 real-world consumer websites (e.g., Amazon, Coursera, Booking), which are used to evaluate WebVoyager, a vision-based web agent.
@@
Our proposed benchmark, WorkArena, is designed to complement existing work by specifically focusing on real-world enterprise software applications. It includes a wide range of tasks that collectively encompass several end-to-end workflows typically performed by knowledge workers. Additionally, it poses a series of technical challenges, such as pages with very large document object models (DOMs), non-standard HTML, and complex UI elements, which we outline in § 3.2. This benchmark integrates into BrowserGym, a new environment that we propose for the evaluation of web agents, which aggregates all features proposed in previous work, such as multimodal observations and code-based actions, while being the first to support chat-based agent-user interactions (§ 4.1).
@@
LLM-based Agents:  The scope of our experimental contributions is limited to web agents that rely on language models for reasoning. Recent studies include the seminal work of Nakano et al. (2021) that introduces WebGPT, an agent capable of browsing the web and answering questions via information retrieval. Other works have also explored web agents that receive HTML as input and produce a series of high-level actions such as click, type, select (Deng et al., 2023; Liu et al., 2023a; b; Yao et al., 2023).
@@
Other works have shown that using just text content as input is limited information, and, therefore, have considered multimodal observations, which includes both visual (screenshots of a page) and text content in order to perform a task (Humphreys et al., 2022; He et al., 2024). Instead of directly interacting with a website, recent works have proposed methods that can act on websites using Python-generated code from task-specific instructions (Gur et al., 2023a; b). Our proposed environment, BrowserGym is flexible in that it supports all observations and actions spaces utilized in prior research.
@@ 3 WorkArena – An Enterprise Benchmark
WorkArena consists of a suite of 29 tasks and 18,050 unique instances that cover core interactions with the ServiceNow platform, such as navigating lists, filling forms, searching knowledge bases, utilizing service catalogs, and navigating via menus (see Fig. 1(a)). Collectively, these tasks are representative of a wide array of common operations that employees, like IT, administrative, and white-collar staff, perform on a daily basis.
As a guiding example, consider an IT support agent tasked with onboarding new hires. Each day, this agent logs into the ServiceNow platform. They use the menu to navigate to a list of requests to be fulfilled. They plan their work by filtering the list to extract all requests assigned to them and sorting the list in order of priority. They then process requests by filling out forms to create new user profiles, and use the service catalog to order laptops for them. As we will see, all of the interactions listed above are included in WorkArena, and this is only one of the many user trajectories that the benchmark covers.
@@ 3.1 WorkArena Tasks
In WorkArena, each task is coupled with a natural language goal that provides unambiguous instructions to the agent (examples are illustrated in Fig. 2). The benchmark includes validation functions that offer real-time feedback to agents, identifying errors ranging from minor (e.g., unfilled mandatory fields) to critical (such as pushing invalid data to the database). A unique feature of WorkArena in contrast with prior benchmarks is the inclusion of a cheating function for each task, implemented using Playwright browser automation (Microsoft, 2023). This function, designed to complete the tasks successfully, serves three purposes: (i) it ensures the feasibility of each task, (ii) it serves as the ground truth for agents that have learning capabilities, and (iii) it helps maintain the benchmark’s longevity by making it easier to identify and adjust tasks impacted by future updates to the ServiceNow platform. The WorkArena benchmark runs on ServiceNow developer instances, which are freely available (with limited capacity) from https://developer.servicenow.com, and initialized to a known state with demonstration data, ensuring consistency and reproducibility of benchmark runs.
@@
Forms:  We consider 5 form-based tasks, which each consist in creating a new entry in a given data table. These tasks vary in complexity based on the number of fields that must be filled (from 1 to 26) and on the intricate properties of each form’s UI. For example, some forms require navigating through a set of tabs to expose hidden fields. Others use dynamic auto-completions fields, which require careful handling (see Fig. 3 for an example). In all cases, validation proceeds by querying the database to retrieve entries created by the agent and verifying that their values are as expected. Together, these tasks yield 5,000 instances.
@@
Service catalogs:  The benchmark includes 9 tasks that require navigating a catalog of products and ordering items with given specifications. Such tasks vary in complexity based on the number of item configuration options. In all such tasks, validation is done by querying the database to verify that the order request created by the agent includes the right items in the right amounts, with the expected specifications. Together, these tasks yield 3,550 instances.
@@
A detailed list of all the tasks implemented in WorkArena is available in § A.1 Tab. 7.
@@ 3.2 Challenges: the World Wild Web of Work
Non-standard, exotic DOMs: Second, the ServiceNow platform relies on a complex combination of web technologies to implement its web pages, with nested iFrames, shadow DOMs, and proprietary Javascript APIs and HTML tags that do not necessarily adhere to existing web standards. [^1] https://www.w3.org/standards/ This specificity would require strong out-of-distribution generalization for a UI assistant to successfully solve a task.
Large DOM trees: Third, the HTML Document Object Model (DOM) of rendered web pages in the ServiceNow platform can be prohibitively large even for state-of-the-art language models, with a flat text size that ranges between 40k and 500k tokens, even after a basic cleaning (removing scripts, styles and empty elements). Thus, even a conceptually simple task, such as finding the next element to click on the current page requires long context understanding, which is an active area of research in language models.
@@ 3.3 Availability
WorkArena is open-source and available to the scientific community. [^2] https://github.com/ServiceNow/WorkArena All tasks are executed in real cloud-based ServiceNow instances, which are freely available via ServiceNow’s developer program (developer.servicenow.com).
@@ 4 BrowserGym
Along with the WorkArena benchmark, we introduce BrowserGym, a generic browser environment that facilitates the design of new benchmarks, and provides a solid platform for the evaluation of multi-modal web agents. BrowserGym (Fig. 1(b)) is implemented as an OpenAI Gym environment (Brockman et al., 2016) and follows a Partially-Observable Markov Decision Process (POMDP) paradigm. [^3] https://github.com/ServiceNow/BrowserGym It relies on Chromium and uses the Chrome DevTools Protocol (CDP) (Google, 2023) and the Playwright library (Microsoft, 2023) to interact with the web browser.
@@ 4.1 Capabilities
Augmented DOM attributes: BrowserGym marks every element on the web pages with a unique identifier bid, its center screen coordinates (x,y), its bounding box (left,top,right,bottom), and a visibility flag (visible). These attributes provide a crude summary of the visual rendering of the UI and allow unambiguous interaction with individual elements through their identifiers.
Rich observation space: at each time step, the observation space contains the content of the chat (list of messages), the currently open pages (list of URLs), the error message from the last action if any (stack trace), and a multi-modal view of the active web page: its DOM snapshot (structured object), its accessibility tree or AXTree (structured object) as originally proposed by Zhou et al. (2023), and a viewport screenshot (image). Both DOM and AXTree are enriched with augmented attributes (bid, coords, visible tag) and are obtained through Chrome’s CDP. These structured objects can typically be rendered as text for processing by a language model and can be combined with the screenshot for a vision-language model.
Rich action space: the action space is customizable and includes Python code, which can be restricted to specific set of pre-defined high-level primitives, such as bid-based actions (click(bid), type(bid,text), …), and coord-based actions (mouse_click(x,y), keyboard_type(text), …). Alternatively, the action space be left unrestricted and allow the execution of arbitrary Python code, including the entire Playwright API and giving the web agent maximum flexibility in interacting with the browser. For the complete list of high-level primitives available in Browsergym, refer to § B, Tab. 8.
@@ 4.2 An Ideal Experimental Framework
Flexible agent design:  BrowserGym offers an extensive list of features but does not impose any restriction on how web agents should be implemented. The agent is responsible for using the provided observations or not (DOM, AXTree, screenshot, error message), deciding how to handle the history (past observations and actions), or deciding which action space it should be using (python, bid, coord, coord+bid). As such, with BrowserGym, researchers can easily experiment with new ideas and evaluate and compare a wide variety of web agents on the same set of tasks, such as text-only agents, vision-augmented agents, memory-augmented agents, and so on.
Minimal task design:  BrowserGym reduces the burden of creating new benchmarks to a minimum. Implementing a new task in BrowserGym boils down to implementing four functions: setup(), teardown(), validate() and cheat() (optional). The setup() function is responsible for initializing anything the task needs beforehand, such as creating database entries, navigating to the starting URL, authenticating, etc. Likewise, teardown() is responsible for cleaning up any resource that might have been created during the task’s execution. validate() is responsible for checking if the task’s goal was fulfilled, which can involve operations such as querying a database, validating the URL and the content of the current page, or looking at the messages sent by the agent in the chat. The method returns a reward, an optional user message for the chat, and a done flag indicating the end of the task. Finally, each task can optionally implement a cheat() function, meant to assess the feasibility of the task via a hard-coded Playwright solution.
@@ 5 Experiments
We present a series of empirical experiments to assess the performance of state-of-the-art, general-purpose LLMs at solving work-related web tasks, using WorkArena and BrowserGym. The aim of these experiments is two-fold. First, we situate the level of difficulty of WorkArena by comparing it across baselines and benchmarks. Second, we propose an incremental analysis to quantify the impact of the different features offered in BrowserGym.

### v5

@@ Abstract
We study the use of large language model-based agents for interacting with software via web browsers. Unlike prior work, we focus on measuring the agents’ ability to perform tasks that span the typical daily work of knowledge workers utilizing enterprise software systems. To this end, we propose WorkArena, a remote-hosted benchmark of 33 tasks based on the widely-used ServiceNow platform. We also introduce BrowserGym, an environment for the design and evaluation of such agents, offering a rich set of actions as well as multimodal observations. Our empirical evaluation reveals that while current agents show promise on WorkArena, there remains a considerable gap towards achieving full task automation. Notably, our analysis uncovers a significant performance disparity between open and closed-source LLMs, highlighting a critical area for future exploration and development in the field.
Machine Learning, ICML
@@
Figure 1: Overview of contributions: (a) WorkArena is a benchmark of 33 web tasks and 19,912 unique instances that cover common ways of interacting with the ServiceNow Platform, a widely-used enterprise software platform. (b) BrowserGym is a Python environment for designing and evaluating web agents, which includes a rich set of actions and multimodal observations (shown here the HTML contents of the page, its accessibility tree, and the raw pixels after browser rendering).
@@ 1 Introduction
  • WorkArena: A realistic benchmark of enterprise-related tasks for web agents comprising 19,912 unique task instances (§ 3, Figure 1a);
  • BrowserGym: A new framework for the development and evaluation of web agents, compatible with previous benchmarks like WebArena (Zhou et al., 2023), MiniWoB (Liu et al., 2018; Shi et al., 2017a) and WebShop (Yao et al., 2022), that offers a richer set of multimodal observations (e.g., screenshot, accessibility tree, screen coordinates), a broader set of actions (e.g., Python code and high-level primitives), and supports chat-based interactions (§ 4, Figure 1b); Surprisingly, these features contribute to bringing our GPT-4 agent at the top of the leaderboard on WebArena, with a score of 25.4%, contrasting with the score of the original paper, 14.4%.
@@ 2 Related Works
Benchmarks for web agents:  Early benchmarks for web agents were based on synthetic web environments where agents were tasked with performing low-level keyboard and mouse actions (Shi et al., 2017b). Notable examples are MiniWoB (Shi et al., 2017a; Liu et al., 2018), which offer a collection of 125 toy web tasks ranging from clicking a specific button to using a basic text editor, and WebShop (Yao et al., 2022), a simulated e-commerce website with shopping tasks that require searching and browsing a catalog of items. More recently, Zhou et al. (2023) introduced WebArena, a collection of 190 tasks based on realistic websites that emulate real-world domains such as e-commerce, social forums, collaborative software development, and content management. WebArena is a notoriously challenging benchmark, with a success rate of 14% for a state-of-the-art web agent based on GPT-4, and 78% for human agents. Deng et al. (2023) proposed Min2Web, a large-scale dataset of 2,000 web interactions from 137 websites curated by human annotators. Similarly, Lù et al. (2024) propose WebLINX, a curated dataset of web interactions composed of 2337 expert demonstrations from 155 different real-world websites. In WebLINX, each task is composed of a turn-based chat dialogue averaging 43 interactions per task. Recently, He et al. (2024) propose 300 information-retrieval tasks, from 15 real-world consumer websites (e.g., Amazon, Coursera, Booking), which are used to evaluate WebVoyager, a vision-based web agent.
Worth mentioning is the related body of work on mobile UI agent benchmarks, which includes the early PixelHelp dataset (Li et al., 2020) and more recent Android in the Wild (Rawles et al., 2023) and Macro Mining (Huang et al., 2024).
@@
Our proposed benchmark, WorkArena, is designed to complement existing work by specifically focusing on real-world enterprise software applications. It includes a wide range of tasks that collectively encompass several end-to-end workflows typically performed by knowledge workers. Additionally, it poses a series of technical challenges, such as pages with very large document object models (DOMs), non-standard HTML, and complex UI elements, which we outline in § 3.2. This benchmark integrates into BrowserGym, a new environment that we propose for the evaluation of web agents, which aggregates all features proposed in previous work, such as multimodal observations and code-based actions while being the first to support chat-based agent-user interactions (§ 4.1).
@@
LLM-based Agents:  The scope of our experimental contributions is limited to web agents that rely on language models for reasoning. Recent studies include the seminal work of Nakano et al. (2021) that introduces WebGPT, an agent capable of browsing the web and answering questions via information retrieval. Other works have also explored web agents that receive HTML as input and produce a series of high-level actions such as click, type, select (Deng et al., 2023; Liu et al., 2023a, b; Yao et al., 2023).
@@
Other works have shown that using only textual information as input is sometimes limiting and have considered multimodal observations that combine visual information (screenshots of a page) with text (Humphreys et al., 2022; He et al., 2024). Instead of directly interacting with a website, recent works have proposed methods that can act on websites using Python-generated code from task-specific instructions (Gur et al., 2023a, b). Our proposed environment, BrowserGym is flexible in that it supports all observations and actions spaces utilized in prior research.
@@ 3 WorkArena – An Enterprise Benchmark
WorkArena consists of a suite of 33 tasks and 19,912 unique instances that cover core interactions with the ServiceNow platform, such as filtering lists, filling forms, searching knowledge bases, utilizing service catalogs, reading dashboards, and navigating the workspace via menus (see Figure 1(a)). Collectively, these tasks are representative of a wide array of common operations that employees, like IT, administrative, and white-collar staff, perform on a daily basis.
As a guiding example, consider an IT support agent tasked with onboarding new hires. Each day, this agent logs into the ServiceNow platform. On the landing page, they see a dashboard showing key statistics of the work assigned to them. Using the workspace’s menu, they navigate to a list of requests to be fulfilled. They then filter the list to extract all requests assigned to them and sort them by priority. Finally, they process each request by filling out forms to create new user profiles and using the service catalog to order laptops for them. As we will see, all of the interactions listed above are included in WorkArena, and this is only one of the many realistic user trajectories that the benchmark covers.
@@ 3.1 WorkArena Tasks
In WorkArena, each task is coupled with a natural language goal that provides instructions to the agent. Each goal is automatically generated from a human-designed template filled with pre-defined values (menu name, field value, item specifications etc.), and explicitely provides all the information required to solve the task (examples in Figure 2). Other key components of tasks include: i) validation functions and ii) oracle functions, a unique feature that distinguishes WorkArena from previous work. The validation functions offer real-time feedback to agents, identifying errors ranging from minor (e.g., unfilled mandatory fields) to critical (such as pushing invalid data to the database). The oracle functions are hand-crafted solutions that automatically complete the tasks using Playwright browser automation (Microsoft, 2023). They serve three purposes: (i) they ensure the feasibility of each task, (ii) they act as ground truth for agents that have learning capabilities, and (iii) they help maintain the benchmark’s longevity by making it easier to identify and adjust tasks impacted by future updates to the ServiceNow platform. Below, we outline the main categories of tasks included in the benchmark.
@@
Forms:  We consider 5 form-based tasks, which each consist of creating a new entry in a given data table. These tasks vary in complexity based on the number of fields that must be filled (from 1 to 26) and on the intricate properties of each form’s UI. For example, some forms require navigating through a set of tabs to expose hidden fields. Others use dynamic auto-completions fields, which require careful handling (see Figure 3 for an example). In all cases, validation proceeds by querying the database to retrieve entries created by the agent and verifying that their values are as expected. Together, these tasks yield 5,000 instances.
@@
Service catalogs:  The benchmark includes 9 tasks that require navigating a catalog of products and ordering items with given specifications. Such tasks vary in complexity based on the number of item configuration options. In all cases, validation is done by querying the database to verify that the order request created by the agent includes the right items in the right amounts, with the expected specifications. Together, these tasks yield 3,550 instances.
Dashboards:  The benchmark includes 4 retrieval tasks that require extracting numerical data from charts and (optionally) performing simple reasoning over them to answer a question. The complexity of such tasks depends on i) whether there are one or many charts on the page and ii) the reasoning to be done (e.g., extract the minimum). In all cases, validation consists of verifying if the agent’s response contains key numbers and labels referring to chart elements. Together, these tasks yield 1,862 instances.
@@
A detailed list of all the tasks implemented in WorkArena is available in § A.1 Table 6.
@@ 3.2 Challenges: the World Wild Web of Work
Non-standard, exotic HTML: Second, the ServiceNow platform relies on a complex combination of web technologies to implement its web pages, with nested iFrames, shadow DOMs, and proprietary Javascript APIs and HTML tags that do not necessarily adhere to existing web standards. [^1] https://www.w3.org/standards/ This specificity would require strong out-of-distribution generalization for a UI assistant to successfully solve a task.
Large HTML: Third, the Document Object Model (DOM) of rendered web pages in the ServiceNow platform can be prohibitively large even for state-of-the-art language models, with a flat HTML text size that ranges between 40k and 500k tokens, even after a basic cleaning (removing scripts, styles and empty elements). Thus, even a conceptually simple task, such as finding the next element to click on the current page requires long context understanding, which is an active area of research in language models.
@@ 3.3 Availability
WorkArena is open-source and designed to be easily extended by the scientific community. [^2] https://github.com/ServiceNow/WorkArena Tasks are executed on Personal Developer Instances, which are real cloud-based instances of the ServiceNow product, accessible for free through their developer program. It is important to note that although our benchmark is built on top of ServiceNow’s platform, it operates independently of any proprietary code.
@@ 4 BrowserGym
Along with the WorkArena benchmark, we introduce BrowserGym, a generic browser environment that facilitates the design of new benchmarks, and provides a solid platform for the evaluation of multi-modal web agents. [^3] https://github.com/ServiceNow/BrowserGym BrowserGym (Figure 1(b)) is implemented as an OpenAI Gym environment (Brockman et al., 2016) and follows a Partially-Observable Markov Decision Process (POMDP) paradigm. [^4] BrowserGym extracts at each time step an observation that is based only on the current view of the page. This emphasizes the need for web agents to implement some kind of memory mechanism, which BrowserGym does not provide. It relies on Chromium and uses the Chrome DevTools Protocol (CDP) (Google, 2023) and the Playwright library (Microsoft, 2023) to interact with the web browser.
@@ 4.1 Capabilities
Augmented attributes: For every element on the current page BrowserGym provides a unique identifier (bid), its bounding box coordinates (left,top,right,bottom), and flags indicating whether it is visible and clickable. These attributes both provide a crude summary of the visual rendering of the UI and allow for unambiguous interaction with individual elements using their identifiers.
Rich observation space: at each time step, the observation space contains the content of the chat (list of messages), the currently open pages (list of URLs), the error message from the last action if any (stack trace), and a multi-modal view of the active web page: its HTML DOM snapshot (structured object), its accessibility tree or AXTree (structured object) as originally proposed by Zhou et al. (2023), and a viewport screenshot (image). Both DOM and AXTree are obtained through Chrome’s CDP, and are enriched with augmented attributes (bid, coords, visible, clickable). These structured objects can typically be rendered as text for processing by a language model and can be combined with the screenshot for a vision-language model.
Rich action space: the action space is customizable and includes Python code, which can be restricted to specific set of pre-defined high-level primitives, such as bid-based actions (click(bid), type(bid,text), …), and coord-based actions (mouse_click(x,y), keyboard_type(text), …). Alternatively, the action space can allow for the execution of arbitrary Python code, including the entire Playwright API which gives the web agent maximum flexibility in interacting with the browser. For the complete list of high-level primitives available in Browsergym, refer to § B, Table 10.
@@ 4.2 An Ideal Experimental Framework
Flexible agent design:  BrowserGym offers an extensive list of features but does not impose any restriction on how web agents should be implemented. The agent is responsible for using the provided observations or not (HTML, AXTree, screenshot, error message), deciding how to handle the history (past observations and actions), or deciding which action space it should be using (python, bid, coord, coord+bid). As such, with BrowserGym, researchers can easily experiment with new ideas and evaluate and compare a wide variety of web agents on the same set of tasks, such as text-only agents, vision-augmented agents, memory-augmented agents, and so on.
Minimal task design:  BrowserGym reduces the burden of creating new benchmarks to a minimum. Implementing a new task in BrowserGym boils down to implementing four functions: setup(), teardown(), validate() and cheat() (an optional oracle function). The setup() function is responsible for initializing anything the task needs beforehand, such as creating database entries, navigating to the starting URL, authenticating, etc. Likewise, teardown() is responsible for cleaning up any resource that might have been created during the task’s execution. validate() is responsible for checking if the task’s goal was fulfilled, which can involve operations such as querying a database, validating the URL and the content of the current page, or looking at the messages sent by the agent in the chat. The method returns a reward, an optional user message for the chat, and a done flag indicating the end of the task. Finally, each task can optionally implement an oracle function, cheat(), that automatically solves the task using a hard-coded Playwright solution. This can notably be used to assess the feasibility of tasks.
@@ 5 Experiments
We present a series of empirical experiments to assess the performance of state-of-the-art, general-purpose LLMs at solving work-related web tasks, using WorkArena and BrowserGym. The aim of these experiments is two-fold. First, we situate the level of difficulty of WorkArena by comparing it across baselines and benchmarks. Second, we propose an ablation study to quantify the impact of the different features offered in BrowserGym. All experiments are run with BrowserGym v0.3.5 and WorkArena v0.3.0.555Code: https://github.com/ServiceNow/AgentLab.
Table 1: Best agent configuration for each LLM, after a random search on MiniWoB and WorkArena.
Type Flag GPT-4o GPT-3.5 Llama3
Agent use_thinking
use_action_history
use_error_history
use_think_history
Obs. Space use_focused_element
use_last_error
coords
extract_visible_tag
extract_clickable_tag
only_visible_elements
Act. Space multi_actions
action_set bid bid bid
individual_examples
long_description